# Is Your Paper Being Reviewed by an LLM? A New Benchmark Dataset and Approach for Detecting AI Text in Peer Review

### v1

@@ 5.1 Fully AI-Written Review Detectability
Table 3: Actual FPR and TPR calculated from the withheld test dataset at varying detection thresholds, which are calibrated using ICLR reviews from our calibration set at different target FPRs. Highest TPRs are in bold.
Reviews generated by open-source LLMs (see Appendix C.2) show somewhat different trends. For both Llama and Qwen reviews, the Binoculars method achieves near-perfect detection performance. While the Anchor method ranks second for Qwen reviews with slightly lower FPRs, it is not as performant on Llama reviews. This is surprising, as Llama reviews appear easier to detect—e.g., nearly twice as many methods achieve TPR@1%FPR > 10% for Llama reviews compared to Qwen, GPT-4o, and Claude reviews. One possibility is that the semantic similarity between the three anchor reviews generated by higher-quality LLMs (GPT-4o, Gemini, and Claude) and Llama reviews is low enough to overlap with that between the anchor reviews and human reviews, potentially blurring the decision boundary. Although the underlying causes of these differences need further investigation, these findings are less central to our study since most LLM users are likely to rely on commercial models due to their ease of use and superior quality.
Method NDCG Minimum Moderate Extensive Maximum Full
Binoculars 0.94 0.14 0.25 0.34 0.56 0.79
Anchor 0.98 0.05 0.09 0.16 0.87 0.94
Table 4: NDCG and the proportion of reviews flagged as AI-written for different levels of editing.
5.2 Detecting AI-Edited Peer Reviews
We also evaluated the ability of our method to correctly rank reviews in ascending order of their likelihood of being AI-generated. The ground truth order is: human-written, minimally-edited, moderately-edited, extensively-edited, maximally-edited, and fully AI-generated reviews. To measure this, we compute the probability that each review type is AI-generated and rank them accordingly. The ranking is compared with the ground truth using the NDCG metric Järvelin and Kekäläinen (2002), with higher values indicating that the predicted ranking is closer to the ground truth ranking. In addition, we analyze the detection rates for each type of review. For each category, we calculate the percentage of reviews identified as AI-generated. Ideally, the detection rate should be lowest for minimally edited reviews and should progressively increase with the extent of AI involvement.
Table 4 compares our method with the best baseline method from our prior experiments (Binoculars). Our anchor embedding method achieves a higher NDCG score compared to the Binoculars approach, indicating that our method ranks each type of review more accurately and closer to the ground truth. Both methods detect increasing percentages of AI-generated text as the extent of editing increases. This suggests that reviews of more extensive editing are more likely to be detected as AI-generated. For Minimum, Moderate, and Extensive types of edited reviews, our method shows a lower AI-generated detection percentage compared to Binoculars. On the other hand, for Maximum and fully AI-generated reviews, our method shows a higher detection rate than Binoculars. This suggests that our approach is better calibrated to detect reviews where AI involvement is predominant.
@@ 6 Analysis
6.1 Human Analysis of Differences Between Human and AI-Written Peer Reviews
@@
To better understand the characteristics which differentiate peer reviews written by humans and LLMs, we conducted a quantitative analysis of 32 reviews authored by humans and GPT-4o for 5 papers submitted to ICLR 2021. Specifically, we read an equal number of human and GPT-4 written reviews for each paper and noted differences in the content between them. A distinguishing characteristic of the analyzed human reviews was that they usually contained details or references to specific sections, tables, figures, or results in the paper. In contrast, peer reviews authored by GPT-4o lacked such specific details, instead focusing on higher-level comments. Another key difference identified in our qualitative analysis was the lack of any specific references to prior or related work in peer reviews generated by GPT-4o. Human authored peer reviews often point out missing references, challenge the novelty of the paper by referencing related work, or suggest specific baselines with references that should be included in the study. In contrast, none of the analyzed GPT-4o reviews contained such specific references to related work. Finally, we found that the vast majority of GPT-4o reviews mentioned highly similar generic criticisms which were not found in human-authored reviews for the same paper. Examples of these issues are provided in Table 13 of Appendix D.1.
@@
Prior work has shown that peer reviews written by GPT-4 and humans have a level of semantic similarity which is comparable to that between different human-authored peer reviews, which has been used to advocate for the usefulness of feedback from GPT-4 in the paper writing process (Liang et al., 2024b). In our qualitative analysis, we found that GPT-4 does indeed generate similar higher-level comments as human reviewers, which could account for this semantic similarity. Despite being generic in nature, we would agree that such feedback could be useful to authors seeking to improve their manuscripts. Nevertheless, we believe that the lack of specificity, detail, and consideration of related work in peer reviews authored by GPT-4 demonstrates that it is not suitable for replacing human domain experts in the peer review process.
6.2 Misalignment Between Human and AI Reviews
@@
In addition to qualitative differences in the content of human and AI-written reviews, we also observe a divergence in numeric scores assigned as part of the review. Figure 3 (Appendix D.2) provides histograms depicting the distribution of score differences for soundness, presentation, contribution, and confidence, which are computed by subtracting scores assigned for each category by human reviewers from those assigned by AI reviewers. AI-written peer reviews were matched with their corresponding human review (aligned by paper ID and overall recommendation) to compute the score differences. Confidence scores range from 1 to 5, while all other categories of scores range from 1 to 4. We focus on reviews from NeurIPS 2022, which were produced prior to the release of ChatGPT. This provides greater confidence that the human-labeled reviews were indeed written by humans, with little to no potential AI influence.
@@
All LLMs produce higher scores than human reviews with a high degree of statistical significance, assessed using a two-sided Wilcoxon signed-rank test (see legend for $p$-values). While the difference between human and AI confidence scores are relatively consistent across all three LLMs, Claude exceeds human scores by the greatest magnitude for soundness, presentation, and contribution. GPT-4o and and Gemini exceed human scores by a similar magnitude for presentation and contribution, while GPT-4o exhibits a greater divergence for soundness scores. Overall these results indicate that AI-written peer reviews are more favorable w.r.t. assigned scores than human-written peer reviews, which raises fairness concerns as scores are highly correlated with acceptance decisions. Our findings are consistent with prior work which has shown that papers reviewed by LLMs have a higher chance of acceptance (Drori and Te’eni, 2024; Latona et al., 2024; Ye et al., 2024).
@@ 7 Conclusion
In this work, we introduced a new large-scale dataset of parallel human and AI-written peer reviews for identical papers submitted to leading AI research conferences. Our evaluations showed that 18 existing open-source approaches for detecting AI text are poorly suited to the challenging problem of identifying AI-generated peer reviews. While high detection rates are possible with existing methods, this comes at the cost of relatively high rates of falsely identifying human-written reviews as containing AI text, which must be minimized in practice. We proposed a new approach which intentionally generates AI-written reviews for a given paper to serve as a basis for comparing semantic similarity to other evaluated reviews, achieving much higher accuracy in identifying AI-written peer reviews while maintaining a low level of false positives. Our results demonstrate the promise of this approach while also motivating the need for further research on methods for detecting unethical applications of LLMs in the peer review process.
Limitations
Our dataset primarily focuses on two conferences, both within the computer science domain. To broaden its applicability and relevance, incorporating additional conferences from diverse research domains would be beneficial. Additionally, our proposed method requires generating reviews for manuscripts, making its use cases more constrained compared to previous approaches. However, in real-world scenarios, when reviewers submit their reviews, they typically accompany the corresponding manuscript. Therefore, our method remains applicable in this setting. Furthermore, our main results are based on evaluations of three commercial LLMs. Given the rapid emergence of new models, conducting comprehensive experiments across all available LLMs is infeasible. In addition, we leverage an open-source platform to conduct the baseline experiments, where baseline performance may vary depending on the choice of surrogate models. However, given the large number of baselines we evaluate, performing an exhaustive search for the optimal surrogate model for each method would be prohibitively expensive. Therefore, we use the default settings. Lastly, we used a fixed set of prompts for different LLMs to generate reviews. In reality, a potential AI-tool reviewer might leverage different prompt and therefore could affect the detection performance.
Ethics Statement
@@
We compare our approach to 18 baseline methods from IGMBT [^2] https://github.com/kinit-sk/IMGTB with its default setting (Spiegel and Macko, 2023), which are categorized into metric-based and pretrained model-based methods. The metric-based methods include Binoculars  Hans et al. (2024), DetectLLM-LLR  Su et al. , DNAGPT  Yang et al. (2023), Entropy  Gehrmann et al. (2019), FastDetectGPT  Bao et al. (2023), GLTR  Gehrmann et al. (2019), LLMDeviation  Wu and Xiang (2023), Loglikelihood  Solaiman et al. (2019), LogRank  Mitchell et al. (2023b), MFD  Wu and Xiang (2023), Rank  Gehrmann et al. (2019), and S5  Spiegel and Macko (2023). The model-based methods include NTNU-D  Sivesind and Winje (2023), ChatGPT-D  Guo et al. (2023), OpenAI-D  Solaiman et al. (2019), OpenAI-D-lrg  Solaiman et al. (2019), RADAR-D  Solaiman et al. (2019), and MAGE-D  Li et al. (2024).
@@ A.1.1 Binoculars
Binoculars  Hans et al. (2024), analyzes text through two perspectives. First, it calculates the log perplexity of the text using an observer LLM. Then, a performer LLM generates next-token predictions, whose perplexity is evaluated by the observer—this metric is termed cross-perplexity. The ratio of perplexity to cross-perplexity serves as a strong indicator for detecting LLM-generated text.
@@ A.1.2 DNAGPT
DNAGPT  Yang et al. (2023) is a training-free detection method designed to identify machine-generated text. Unlike conventional approaches that rely on training models, DNAGPT uses Divergent N-Gram Analysis (DNA) to detect discrepancies in text origin. The method works by truncating a given text at the midpoint and using the preceding portion as input to an LLM to regenerate the missing section. By comparing the regenerated text with the original through N-gram analysis (black-box) or probability divergence (white-box), DNAGPT reveals distributional differences between human and machine-written text, offering a flexible and explainable detection strategy.

### v2

@@ 4.1 Fully AI-Written Review Detectability
Table 2 provides these results separately for the detection of GPT-4o, Gemini, and Claude reviews (Llama and Qwen results are provided in Appendix C.3). Other baseline methods that failed achieve a TPR of at least 1% at a target FPR of 1% (i.e., TPR $<1\%$ at target FPR $=1\%$) are omitted.
The results in Table 2 show that among the 18 baselines, Binoculars generally achieves the best performance overall. For example, its TPR reaches 45–85% at a 1% target FPR, while other models reach only up to 19% TPR. However, even with the best-performing Binoculars, performance drops significantly at more stringent FPR levels—for instance, it achieves only 17% TPR at 0.1% FPR for GPT-4o reviews. These results highlight a key challenge in AI-generated peer review detection: the difficulty of robustly identifying AI-written text at the level of individual peer reviews. At the same time, they point to the need for new methods that can improve accuracy and robustness under low-FPR constraints. To this end, we explore a purpose-built detection method (Anchor) in Section 5.1, which leverages additional context specific to the peer review detection task.
@@ 5 Analysis
5.1 Can manuscript context improve detection of AI-generated peer reviews?
Unlike general AI text detection scenarios, peer review provides additional contextual information for this problem: the manuscript being reviewed. With access to metadata connecting reviews to their source manuscripts (as in platforms like OpenReview.net), we investigate whether leveraging the manuscript can improve the detection of AI-generated peer reviews.
To test this idea, we introduce a method which utilizes the manuscript by comparing the semantic similarity between a test reviews (TR) and a synthetic "Anchor Review" (AR) generated for the same manuscript. The AR can be generated by any LLM. We use a simple, generic prompt (Appendix E.2) to generate the AR without prior knowledge of the user prompts (i.e., the AR prompt differs from those used to create reviews in the testing dataset). Once an AR is generated for a given paper (Eq.1) and a testing review (TR) is provided, we obtain their embeddings using a text embedding model (EM, Eqs.2 and 3). The semantic similarity between the embeddings of the AR and TR is then computed using the cosine similarity function (Eq.4). Finally, this similarity score is compared against a learned threshold ($\theta$): if the score exceeds the threshold, the review is classified as AI-generated; otherwise, it is not (Eq. 5). Putting everything together, the method is formalized as:
$\text{AR}=\text{LLM}(\text{paper},\text{Prompt}_{\text{AR}})$ (1)
[…]
In our study, we use OpenAI’s embedding model (text-embedding-003-small). The threshold $\theta$ is learned from the calibration set. Specifically, for each review in the calibration data, we apply the steps outlined in Eqs. 1 to 4. To handle cases where the source LLM is unknown, we generate multiple anchor reviews using different LLMs and apply a voting scheme: if any anchor yields a positive detection, the review is labeled as AI-generated (see Appendix D.1 for details).
Compared to the 18 baseline models, our anchor embedding approach consistently achieves the highest AUC for AI peer review (Figure 2) and outperforms other baseline methods overall (Table 2). The performance gap is especially notable for GPT-4o, which is generally harder to detect; for example, at a target FPR of 0.1%, the anchor embedding method achieves a TPR of 63.5% for GPT-4o reviews, compared to 17.1% for the next-best method (Binoculars), which is an absolute improvement of 46.4%. This indicates that the anchor embedding approach is particularly effective for the most challenging review samples. On Gemini reviews, where baseline methods already perform well, the anchor embedding method performs similarly to Binoculars. These results suggest that leveraging the context of the submitted manuscript through anchor embeddings offers substantial benefits for high-risk, low-FPR detection tasks, especially when reviews are written by the most advanced LLMs.
5.2 Is AI peer review detection robust to prompt variations?
A potential limitation in constructing AI-generated datasets is the reliance on a fixed prompt, which may constrain stylistic diversity or lead to overfitting of downstream detection models. Moreover, such fixed prompting may not reflect real-world use cases, where different users are likely to employ varied and personalized prompts when generating reviews. While our primary prompt is designed to be semantically grounded—using the paper content and a human reviewer’s overall score to guide generation—it is important to evaluate whether this design leads to narrow outputs or brittle detection performance. To assess the impact of prompt variation, we conducted a prompt sensitivity analysis using an alternative prompting strategy that simulates diverse reviewer archetypes (e.g., "balanced," "conservative," "innovative," and "nitpicky"; see Appendix E.4 for the exact prompts we used). For this analysis, we generated 1,921 GPT-4o reviews for papers corresponding to the ICLR 2021 test set.
Cross-Prompt Detection Robustness
We evaluated whether prompt-induced shifts in review style affect the robustness of AI text detectors. Specifically, we applied detection models calibrated on the main (i.e., score-aligned) prompt and evaluated them on reviews generated using the archetype-based prompt. Table S10 reports the TPRs at multiple target FPRs for a range of baseline detection methods. Across all settings, we observe that TPRs under cross-prompt testing remain largely consistent with those under in-prompt testing, with minimal change in detection performance. This result suggests that, despite changes in prompt framing, the core distributional properties of the AI-generated reviews are sufficiently stable for detection models to generalize. Overall, these findings indicate that the utility of our dataset is not overly sensitive to prompt selection.
Embedding-Based Style Consistency
We further investigated prompt-induced variation using sentence-level embedding analysis. Figure S2 visualizes a t-SNE projection of review embeddings from two groups: (i) AI-generated reviews using the primary score-aligned prompt and (ii) AI-generated reviews using the archetype prompt. No meaningful separation emerged between these two review types, suggesting they share broadly similar linguistic and semantic characteristics. This finding reinforces the robustness of our prompting strategy and implies that moderate prompt variation does not lead to drastic stylistic shifts in model outputs.
5.3 Can detection models distinguish human reviews edited by LLMs?
LLMs are widely used for writing assistance tasks such as grammar correction and fluency enhancement [52, 53], especially benefiting non-native speakers. However, excessive reliance on LLMs can lead to substantial rewriting, blurring the line between human- and AI-authored reviews. To simulate this scenario, we took human-written peer reviews from the ICLR 2021 GPT-4o test subset and generated AI-edited versions at four increasing levels of editing: Minimum, Moderate, Extensive, and Maximum (see Appendix E.3 for prompts). To validate that the edits are semantically distinct, we measured cosine similarity between the original and edited reviews using a sentence embedding model. As shown in Table S12, similarity scores decrease as the level of editing increases.
We evaluate detection models on two fronts: (i) their ability to flag edited reviews as AI-generated, and (ii) their ability to rank the reviews by degree of AI involvement using Normalized Discounted Cumulative Gain (NDCG) [54]. Table 3 compares our anchor embedding method with Binoculars, the strongest-performing baseline from previous experiments (as shown in Table 2). Both models achieve high NDCG scores (0.90 for Anchor, 0.86 for Binoculars), indicating that they can generally rank reviews according to the degree of AI editing. However, their flagging behavior differs. At the 0.1% FPR threshold, both models rarely flag minimally to extensively edited reviews as AI-generated ($\leq$2.5%), but diverge sharply at the Maximum level: Anchor flags 60.8% of these reviews, compared to just 9.2% for Binoculars. This trend persists at the 1% FPR threshold, with generally higher flagging rates across all levels. These results suggest that the Anchor method is more sensitive to high levels of AI involvement, whereas Binoculars is more conservative at the upper end.
Overall, these results suggest that high-performing detection models can distinguish varying levels of AI editing in human-written reviews. However, lightly edited texts remain difficult to detect, highlighting a challenge for future work on identifying hybrid human–AI content.
Table 3: NDCG and the proportion of reviews flagged as AI-written for different levels of editing. Flagging rates are shown for two different thresholds (0.1% and 1% target FPRs), adapted from those used in Table 2.
Method NDCG (Threshold: 0.1%FPR) (Threshold: 1%FPR)
Minimum Moderate Extensive Maximum Minimum Moderate Extensive Maximum
Anchor 0.90 0.4 0.7 1.9 60.8 2.6 4.9 9.7 82.3
Binoculars 0.86 0.6 1.4 2.5 9.2 2.7 5.3 9.4 26.4
5.4 How do human-written and AI-generated peer reviews differ?
@@
To better understand the characteristics which differentiate peer reviews written by humans and LLMs, we conducted a quantitative analysis of 32 reviews authored by humans and GPT-4o for 5 papers submitted to ICLR 2021. Specifically, we read an equal number of human and GPT-4o written reviews for each paper and noted differences in the content between them. A distinguishing characteristic of the analyzed human reviews was that they usually contained details or references to specific sections, tables, figures, or results in the paper. In contrast, peer reviews authored by GPT-4o lacked such specific details, instead focusing on higher-level comments. Another key difference identified in our qualitative analysis was the lack of any specific references to prior or related work in peer reviews generated by GPT-4o. Human-authored peer reviews often point out missing references, challenge the novelty of the paper by referencing related work, or suggest specific baselines with references that should be included in the study. In contrast, none of the analyzed GPT-4o reviews contained such specific references to related work. Finally, we found that the vast majority of GPT-4o reviews mentioned highly similar generic criticisms which were not found in human-authored reviews for the same paper. Examples of these issues are provided in Table S11 in Appendix D.3.
@@
Prior work has shown that peer reviews written by GPT-4 and humans have a level of semantic similarity which is comparable to that between different human-authored peer reviews, which has been used to advocate for the usefulness of feedback from GPT-4 in the paper writing process [37]. In our qualitative analysis, we found that GPT-4 does indeed generate similar higher-level comments as human reviewers, which could account for this semantic similarity. Despite being generic in nature, we would agree that such feedback could be useful to authors seeking to improve their manuscripts. Nevertheless, we believe that the lack of specificity, detail, and consideration of related work in peer reviews authored by GPT-4 demonstrates that it is not suitable for replacing human domain experts in the peer review process.
5.5 Do AI-generated reviews assign higher scores than human reviews?
@@
In addition to qualitative differences in the content of human and AI-written reviews, we also observe a divergence in numeric scores assigned as part of the review. Figure S3 in Appendix D.4 provides histograms depicting the distribution of score differences for soundness, presentation, contribution, and confidence, which are computed by subtracting scores assigned for each category by human reviewers from those assigned by AI reviewers. AI-written peer reviews were matched with their corresponding human review (aligned by paper ID and overall recommendation) to compute the score differences. Confidence scores range from 1 to 5, while all other categories of scores range from 1 to 4. In the following discussion, We focus on reviews from NeurIPS 2022, which were produced prior to the release of ChatGPT. This provides greater confidence that the human-labeled reviews were indeed written by humans, with little to no potential AI influence.
@@
All LLMs produce higher scores than human reviews with a high degree of statistical significance, assessed using a two-sided Wilcoxon signed-rank test (see legend for $p$-values in Figure S3). While the difference between human and AI confidence scores are relatively consistent across all three LLMs, Claude exceeds human scores by the greatest magnitude for soundness, presentation, and contribution. GPT-4o and and Gemini exceed human scores by a similar magnitude for presentation and contribution, while GPT-4o exhibits a greater divergence for soundness scores. Overall these results indicate that AI-generated peer reviews are more favorable w.r.t. assigned scores than human-written peer reviews, which raises fairness concerns as scores are highly correlated with acceptance decisions. Our findings are consistent with prior work which has shown that papers reviewed by LLMs have a higher chance of acceptance [10, 6, 11].
6 Limitations
Our dataset primarily focuses on two conferences, both within the computer science domain. To broaden its applicability and relevance, incorporating additional conferences from diverse research areas would be beneficial. While we designed our prompting strategy to encourage stylistic diversity, prompt choice still influences generation, and real-world use cases may involve a wider range of prompting styles than those tested. We conducted a prompt sensitivity analysis (Section 5.2) to evaluate the robustness of our dataset under prompt variation, though broader coverage remains an open direction. Our study also includes detection scenarios where LLMs revise or extend human-written reviews ("AI-edits-human"), but does not simulate the reverse case—where a human revises an AI-generated draft—due to the difficulty of sourcing domain experts to perform such edits at scale. We consider this an important direction for future work. Lastly, our main results are based on evaluations of three commercial LLMs. Given the rapid emergence of new models, conducting comprehensive experiments across all available LLMs is infeasible. In addition, we leverage an open-source platform to run baseline experiments, where performance may vary depending on the choice of surrogate models. However, given the large number of baselines we evaluate, performing an exhaustive search for the optimal surrogate model for each method would be prohibitively expensive. Therefore, we use the default settings.
While our anchor embedding method shows promising results, it is not without limitations. It is task-specific by design and is therefore unsuitable as a general-purpose AI text detection model. Additionally, its reliance on commercial LLM APIs may introduce challenges w.r.t. computational cost and scalability. The method’s performance can also be sensitive to the choice of anchor LLMs. This is not specific to Anchor but reflects a general challenge in black-box detection settings where the source LLM is unknown. Acknowledging these limitations, our findings nonetheless highlight how leveraging manuscript context as auxiliary information can significantly improve the accuracy and robustness of AI-generated peer review detection, especially under low-FPR constraints.
@@ 7 Conclusion
In this work, we introduced a new large-scale dataset of parallel human-written and AI-generated peer reviews for identical papers submitted to leading AI research conferences. Our evaluations show that existing open-source methods for AI text detection struggle in the peer review setting, where high detection rates often come at the cost of falsely flagging human-written reviews—an outcome that must be minimized in practice. We demonstrate that leveraging manuscript context is a promising strategy for improving detection accuracy while maintaining a low false positive rate. In addition, AI-generated reviews tend to be less specific and less grounded in the manuscript than human-written ones. We also find that AI-generated reviews consistently assign higher scores, raising fairness concerns in score-driven decision-making processes. We hope our results motivate further research on responsible detection of AI-generated content in scientific review workflows, and that our dataset provides a valuable resource for advancing this goal.
@@
We compare our approach to 18 baseline methods from IGMBT [^3] https://github.com/kinit-sk/IMGTB (released under MIT license) with its default setting [34], which are categorized into metric-based and pretrained model-based methods. The metric-based methods include Binoculars  [55], DetectLLM-LLR  [36], DNAGPT  [49], Entropy  [31], FastDetectGPT  [56], GLTR  [31], LLMDeviation  [57], Loglikelihood  [27], LogRank  [58], MFD  [57], Rank  [31], and S5  [34]. The model-based methods include NTNU-D  [59], ChatGPT-D  [20], OpenAI-D  [27], OpenAI-D-lrg  [27], RADAR-D  [27], and MAGE-D  [60].
@@ A.1.1 Binoculars
Binoculars [55] analyzes text through two perspectives. First, it calculates the log perplexity of the text using an observer LLM. Then, a performer LLM generates next-token predictions, whose perplexity is evaluated by the observer—this metric is termed cross-perplexity. The ratio of perplexity to cross-perplexity serves as an indicator for detecting LLM-generated text.
@@ A.1.2 DNAGPT
DNAGPT [49] is a training-free detection method designed to identify machine-generated text. Unlike conventional approaches that rely on training models, DNAGPT uses Divergent N-Gram Analysis (DNA) to detect discrepancies in text origin. The method works by truncating a given text at the midpoint and using the preceding portion as input to an LLM to regenerate the missing section. By comparing the regenerated text with the original through N-gram analysis (black-box) or probability divergence (white-box), DNAGPT reveals distributional differences between human and machine-written text, offering a flexible and explainable detection strategy.